# In search of the electron-phonon contribution to total energy

### v1

@@ Abstract
The total energy is a fundamental characteristic of solids, molecules, and nanostructures. In most first-principles calculations of the total energy, the nuclear kinetic operator is decoupled from the many-body electronic Hamiltonian and the dynamics of the nuclei is reintroduced afterwards. This two-step procedure introduced by Born and Oppenheimer (BO) is approximate. Energies beyond the electronic and vibrational (or phononic) main contributions might be relevant when small energy differences are important, such as when predicting stable polymorphs or describing magnetic energy landscape. We clarify the different flavors of BO decoupling and give an exact formulation for the total energy in the basis of BO electronic wavefunctions. Then, we list contributions, beyond the main ones, that appear in a perturbative expansion in powers of $M_{0}^{-1/4}$, where $M_{0}$ is a typical nuclear mass, up to sixth order. Some of these might be grouped and denoted the electron-phonon contribution to total energy, $E^{\textrm{elph}}$, that first appears at fourth order. The electronic inertial mass contributes at sixth order. We clarify that the sum of the Allen-Heine-Cardona zero-point renormalization of eigenvalues over occupied states is not the electron-phonon contribution to the total energy but a part of the phononic contribution. The computation of the lowest-order $E^{\textrm{elph}}$ is implemented and shown to be small but non-negligible (3.8 meV per atom) in the case of diamond and its hexagonal polymorph. We also estimate the electronic inertial mass contribution and confirm the size-consistency of all computed terms.
@@ I Introduction
The total energy is a cornerstone of modern solid-state physics and quantum chemistry. It governs material stability [Jones2015], phase transition [Wentzcovitch1991], crystal structure predictions [Payne1992], or defect formation energy [Freysoldt2014]. Density functional theory (DFT) is often the method of choice when computing the total energy of solids [Hohenberg1964, Kohn1965]. Indeed, if we had access to the exact exchange-correlation functional, DFT would give the exact electronic total energy of crystals. However, DFT rests on the Born-Oppenheimer (BO) decoupling of nuclear and electronic degrees of freedom [Born1927, Born1954]. Often, the first-principles BO nuclear dynamic contribution is added to the electronic energy. Phononic (or vibrational) degrees of freedom are taken into account through Bose-Einstein statistics in the case of harmonic solids [Baroni2001], or through the quasiharmonic approximation to obtain thermal expansion [Rignanese1996, Rostami2024, Rostami2025] or using effective interatomic force constants to include anharmonic (phonon-phonon) corrections [Hellman2011, Monacelli2021, Knoop2024] However, the BO decoupling leaves away numerous contributions to the total energy, beyond the dominant electronic (static nucleus) energy and the phononic harmonic, quasi-harmonic, or anharmonic contributions. So, one can wonder whether electron-phonon contributions to the total energy exist, how to characterize them, and how to compute them.
@@
As a justification for their approach, Born and Oppenheimer [Born1927] relied on a perturbation expansion, using $M_{0}^{-1/4}$ as a small parameter, where $M_{0}$ is a typical nuclear mass. In the equilibrium BO geometry, the total energy is obtained for frozen nuclei, so they do not depend on $M_{0}$, while the first-order contribution vanishes. The second-order contribution, scaling as $M_{0}^{-1/2}$, is the phononic contribution. The third-order correction also vanishes. Additional corrections come at the fourth order, with $M_{0}^{-1}$ scaling. Among these, two contributions are of the anharmonic phonon-phonon type, and can nowadays be included in first-principles calculations of total energy [Hellman2011, Monacelli2021, Knoop2024].
@@
In this work, we observe that the other BO expansion fourth-order correction can be computed approximately (within DFT) on the basis of the electron-phonon matrix elements. Thanks to recent advances in efficient calculations of electron-phonon interaction [Giustino2017], mostly relying on density functional perturbation theory [Baroni1987, Gonze1997, Gonze1997b, Baroni2001, Gonze2005a, Gonze2024a], electron-phonon matrix elements are readily available when computing electronic [Ponce2020] or thermal transport [Allen1993, Lindsay2016] properties, superconductivity [Margine2013], polaron formation [Sio2019, Vasilchenko2025], or the temperature-dependence of the band gap and its zero-point renormalization [Ponce2015]. Several terms of even higher order in the small parameter $M_{0}^{-1/4}$ fall into that category. All could be grouped as “electron-phonon contributions to the total energy”. It is expected that such contributions are small compared to electronic and phononic contributions but they might be crucial for the study of defects and surfaces [Freysoldt2018], magnetic configurations [Cerny2003], chemical reactions and bonding [Xu2020], or accurate phase transition [Bouchet2019] where small energy differences matter.
@@
Given the recent advances in the computation of electron-phonon matrix elements, several recent works also discussed electron-phonon contributions to total energy. In 2020, P. B. Allen [Allen2020] suggested that this contribution could be calculated using ingredients available when electron-phonon band-gap renormalization calculations are performed. His approach is based on the summation over occupied states of partial Fan-Migdal and Debye-Waller self-energies [Allen1976]. Recently, S. Paul et al. [Paul2023] have performed first-principles calculations based on this idea. They argued that such formulation violates size consistency, in which the total energy would depend on the size of the unit cell. In the present work, we show that the formulation of Allen is well defined and size-consistent. Crucially, we show that the approach proposed by Allen does not deliver the fourth-order $M_{0}^{-1}$ electron-phonon contribution term but is already part of the second-order phononic $M_{0}^{-1/2}$ contribution to the total energy, and should not be added to total energy calculations.
@@
We first present in Section II the BO Hamiltonian, and approximations based on BO wavefunctions. We also derive an expression for the exact total energy that includes contributions beyond the BO wavefunctions. We then perform a detailed mass-scaling analysis of all the terms entering the total energy in Section III. From this mass scaling analysis, we identify which terms are phononic contributions, Section IV, while the remaining ones are detailed in Section V. In particular, we report, for the first time explicitly, the lowest-order electron-phonon contribution to the total energy. In Appendix A, the BO perturbation expansion is derived and expressed in normal mode coordinates. We then show in Section VI the extension to solids with practical formula for implementation in first-principles software. In Section VII, we show that Allen’s energy is not the electron-phonon contribution to the total energy but a part of the phononic contribution to the total energy. We then recast in Section VIII Allen’s formula into the standard Allen-Heine-Cardona (AHC) [Allen1976, Allen1981] theory for band renormalization thanks to an identically null term. In Section IX, we present a detailed analysis of the energy terms at the phonon zone-center, comparing a frozen-phonon approach and perturbation theory. We find excellent agreement for all terms, below 0.5% differences. In Section X, we show size consistency by computing the diamond band renormalization using the Abinit code [Gonze2016, Gonze2020, Romero2020, Verstraete2025] and an interpolation of the perturbed potential which includes dipole and quadrupole contributions [Brunin2020, Brunin2020a]. We show that the same results are obtained between the primitive and a supercell of diamond when no crystal symmetries are used.
@@
Finally, in Section XI we compute the (fourth order) electron-phonon contribution to the total energy and approximate the effect of electron mass rescaling (sixth order), for two allotropes of carbon: the stable face-centered cubic diamond and the high-pressure metastable hexagonal diamond called lonsdaleite which is typically found in meteorites. For both materials, we report their converged total energy with momentum integration. We compute the electron-phonon forces and the new equilibrium lattice parameters of diamond and internal parameter of lonsdaleite. At 0 K, we find that diamond is 52.7 meV/(2 atoms) more stable than lonsdaleite.
@@ II Born-Oppenheimer-based approximations
We first consider finite systems and extend the theory to periodic crystals in Sec. VI. The total Hamiltonian of a system composed of electrons and nuclei can be expressed as [Born1927, Born1996]
@@
where $\{{\mathbf{R}}\}$ denotes the set of nuclear positions, $\hat{T}^{\textrm{N}}$ is the nuclear kinetic operator, $\hat{P}_{\kappa\alpha}\equiv-i\partial/\partial R_{\kappa\alpha}$ is the canonical momentum operator for nuclei, with $M_{\kappa}$ the mass of nucleus $\kappa$ and $\alpha$ an index for cartesian axes. Electronic coordinates are implicit, unlike the nuclear ones. The global Hamiltonian in Eq. (1) is written in calligraphic style as well as the forthcoming associated total energies, that are independent of $\{{\mathbf{R}}\}$. Hartree atomic units are used throughout.
@@
The denominations “Born-Oppenheimer Hamiltonian”, “Born-Oppenheimer wavefunction” and “Born-Oppenheimer eigenenergy” for quantities in Eqs. (3) and (6) are well-established and consistent in the literature, but there is some confusion about the meaning of “Born-Oppenheimer approximation” and “adiabatic approximation” [Ballhausen1972, Nafie1983, Hagedorn1986, Tully2000, Scherrer2017, Cohen2025]. In many publications, these are used interchangeably while in other they have different meanings. For example, in their 1972 publication, Ballhausen and Hansen [Ballhausen1972] discuss various adiabatic approximations including “Born-Oppenheimer”, “Born-Huang”, “Herzberg-Teller”, and “crude”, based on wavefunctions of the type
@@
Figure 1: Example of coordinate diagram for the BO energy hypersurfaces $E^{\rm BO}_{j}\{{\mathbf{R}}\}$ of the electronic ground state ($j=0$) and first- and second-excited states ($j=1$ and $j=2$). Eigenenergies of the total Hamiltonian are also shown, labelled with the $i$ index. The ground state of the total Hamiltonian has $i=0$ label. Usually, the corresponding wavefunctions will be made predominantly of the BO wavefunctions of one of the BO hypersurfaces, but there will be hybridized states. The chosen reference configuration is denoted by $\{\mathbf{R}^{\rm ref}\}$. In this example, we present a specific case where the $i=6$ and $i=7$ energy levels are degenerate, being hybrids from the $j=0$ and $j=1$ BO wavefunctions, and where the $j=0$ BO hypersurface has three minima, one global one and two local ones, that might all be chosen to be the reference state $\{\mathbf{R}_{j=0}^{0}\}$ in Sec. III. The $j=2$ one has one global minimum.
Assuming the smooth restricted BO version of Eq. (7), a first possible approximation consists in considering $E^{\rm BO}_{j}\{{\mathbf{R}}\}$ as an effective scalar potential for the nuclear dynamics:
@@
A second possible approximation, which we call the “optimal approximation”, is based on choosing $\chi_{ij}\{{\mathbf{R}}\}$ such that the expectation value of the Hamiltonian, Eq. (1), is stationary (minimal in the ground-state case). In that case, the nuclear wave function fulfills the following equation:
@@
where the projector on the $j$-BO manifold and its orthogonal are defined as
$\hat{\cal{P}}^{\textrm{BO}}_{j}\equiv$ $|\displaystyle\int\mathrm{d}\{{\mathbf{R}}\}\ket{\phi_{j}\{{\mathbf{R}}\}}\bra{\phi_{j}\{{\mathbf{R}}\}}$ (18)
@@
The presence of the projector in Eq. (17) removes the phase arbitrariness that is present in Eq. (12).
In the literature dealing with phonon chirality [Bonini2023], a simplified version of Eq. (13) is often encountered. We refer to it as an “hybrid approximation” since it involves a scalar and a vector potential but does not consider the Born term:
$\Big(\hat{T}^{\textrm{vec}}_{j}\{{\mathbf{R}}\}\!+\!E^{\rm BO}_{j}\{{\mathbf{R}}\}\Big)\chi_{ij}^{\rm hyb1}\{{\mathbf{R}}\}={\cal{E}}_{ij}^{\rm hyb1}\chi_{ij}^{\rm hyb1}\{{\mathbf{R}}\}.$ (20)
Alternatively, before the work of Mead and Truhlar, another hybrid was also present [Born1954], in which the Born term is retained, and the vector potential is ignored:
$\Big(\hat{T}^{\textrm{N}}+E^{\rm opt}_{j}\{{\mathbf{R}}\}\Big)\chi_{ij}^{\rm hyb2}\{{\mathbf{R}}\}={\cal{E}}_{ij}^{\rm hyb2}\chi_{ij}^{\rm hyb2}\{{\mathbf{R}}\}.$ (21)
All four approximations, Eqs. (9), (13), (20) and (21), to the exact Schrödinger equation, Eq. (1), have been referred to as “adiabatic approximation” and/or as “Born-Oppenheimer approximation” in the literature, with the following approximate total wavefunctions:
@@
[…]
Different approximate treatments beyond these four approximations have also been explored by Nafie [Nafie1983], although the vector potential was not considered in that work.
@@
with the term between parentheses being a pseudo-inverse, limited to that subspace. Although Green’s function might not be defined for specific real $z$ values above the ground-state energy $z=\mathcal{E}_{0}$, the restriction to the space orthogonal to the BO-wavefunctions removes this problem up to $z$ being equal to the lowest excited-state BO eigenenergy. We also assume that $z=\mathcal{E}_{i}$ is a value for which the projected Green’s function $\hat{\mathcal{G}}_{\perp j}(z,\{{\mathbf{R}}\})$ exists.
@@
is an operator similar to self-energy operators found in the many-body perturbation theory context [Mahan1990].
Importantly, Eq. (30) is an exact version of the approximate Eqs. (9), (13), (20), (21) and focuses only on nuclear dynamics where all the electronic degrees of freedom are condensed in the self-energy operator. The Green’s function of the full Hamiltonian fulfills a Dyson-type equation:

### v2

@@ Abstract
The total energy is a fundamental characteristic of solids, molecules, and nanostructures. In most first-principles calculations of the total energy, the nuclear kinetic operator is decoupled from the many-body electronic Hamiltonian and nuclear potential, and the dynamics of the nuclei is reintroduced afterward. This two-step procedure introduced by Born and Oppenheimer (BO) is approximate. Energies beyond the electronic and vibrational (or phononic) main contributions might be relevant when small energy differences are important, such as when predicting stable polymorphs or describing magnetic energy landscape. We clarify the different flavors of BO decoupling and give an exact formulation for the total energy in the basis of BO electronic wavefunctions. Then, we list contributions, beyond the main ones, that appear in a perturbative expansion in powers of $M_{0}^{-1/4}$, where $M_{0}$ is a typical nuclear mass, up to sixth order. Some of these might be grouped and denoted the electron-phonon contribution to total energy, $E^{\textrm{elph}}$, that first appears at fourth order. The electronic inertial mass contributes at sixth order. We clarify that the sum of the Allen-Heine-Cardona zero-point renormalization of eigenvalues over occupied states is not the electron-phonon contribution to the total energy but a part of the phononic contribution. The computation of the lowest-order $E^{\textrm{elph}}$ is implemented and shown to be small but non-negligible (3.8 meV per atom) in the case of diamond and its hexagonal polymorph. We also estimate the electronic inertial mass contribution and the quasi-harmonic one. We confirm the size consistency of all computed terms.
@@ I Introduction
The total energy is a cornerstone of modern solid-state physics and quantum chemistry. It governs material stability [40], phase transition [91], crystal structure predictions [56], or defect formation energy [23]. Density functional theory (DFT) is often the method of choice when computing the total energy of solids [37, 42]. If the exact exchange-correlation functional was known, then for any system of nuclei and electrons, DFT would give the exact total electronic energy at any chosen set of fixed nuclear positions $\{\mathbf{R}\}$. An open question is how to get the exact total energy (i.e. electronic plus nuclear vibrational) for the ground state of such a system. Indeed, DFT rests on the Born-Oppenheimer (BO) decoupling of nuclear and electronic degrees of freedom [13, 12]. Often, the first-principles BO nuclear contribution is added to the electronic energy. Phononic (or vibrational) degrees of freedom are taken into account through Bose-Einstein statistics in the case of harmonic solids [8], or through the quasiharmonic approximation to obtain thermal expansion [70, 72, 73] or using effective interatomic force constants to include anharmonic (phonon-phonon) corrections [36, 53, 41]. However, the BO decoupling leaves away numerous contributions to the total energy beyond the dominant electronic (static nucleus) energy and the phononic harmonic, quasi-harmonic, or anharmonic contributions. So, one can wonder whether electron-phonon contributions to the total energy exist, how to characterize them, and how to compute them.
@@
As a justification for their approach, Born and Oppenheimer [13] relied on a perturbation expansion, using $M_{0}^{-1/4}$ as a small parameter, where $M_{0}$ is a typical nuclear mass. In the equilibrium BO geometry, the total energy is obtained for frozen nuclei, so it does not depend on $M_{0}$, while the first-order contribution vanishes. The second-order contribution, scaling as $M_{0}^{-1/2}$, is the phononic contribution. The third-order correction also vanishes. Additional corrections come at the fourth order, with $M_{0}^{-1}$ scaling. Among these, two contributions are of the anharmonic phonon-phonon type, and can nowadays be included in first-principles calculations of total energy [36, 53, 41].
@@
In this work, we observe that the other fourth-order BO expansion correction can be computed approximately (within DFT) on the basis of the electron-phonon matrix elements. Thanks to recent advances in efficient calculations of electron-phonon interaction [26], mostly relying on density functional perturbation theory [9, 33, 30, 8, 31, 32], electron-phonon matrix elements are readily available when computing electronic transport [62, 45] or thermal transport [4, 46] properties, superconductivity [49], polaron formation [82, 88], or the temperature-dependence of the band gap and its zero-point renormalization [60, 63]. Several terms of even higher order in the small parameter $M_{0}^{-1/4}$ fall into that category. All could be grouped as “electron-phonon contributions to the total energy”. It is expected that such contributions are small compared to electronic and phononic contributions, but they might be crucial for the study of defects and surfaces [24], magnetic configurations [20], chemical reactions and bonding [92], or accurate phase transition [15] where small energy differences matter.
@@
Given the recent advances in the computation of electron-phonon matrix elements, several recent works also discussed electron-phonon contributions to total energy. In 2020, P. B. Allen [5] suggested that this contribution could be calculated using ingredients available when electron-phonon band-gap renormalization calculations are performed. His approach is based on the summation over occupied states of partial Fan-Migdal and Debye-Waller self-energies [2]. Recently, S. Paul et al. [55] have performed first-principles calculations based on this idea. They argued that such formulation violates size consistency, in which the total energy would depend on the size of the unit cell. In the present work, we show that the formulation of Allen is well defined and size-consistent. Crucially, we show that the approach proposed by Allen does not deliver the fourth-order $M_{0}^{-1}$ electron-phonon contribution term but is already part of the second-order phononic $M_{0}^{-1/2}$ contribution to the total energy and should not be added to total energy calculations.
@@
We first present in Section II the BO Hamiltonian, and approximations based on BO wavefunctions. We also derive an expression for the exact total energy that includes contributions beyond the BO wavefunctions. We then perform a detailed mass-scaling analysis of all the terms entering the total energy in Section III. From this mass scaling analysis, we identify which terms are phononic contributions, Section IV, while the remaining ones are detailed in Section V. In particular, we report, for the first time explicitly, the lowest-order electron-phonon contribution to the total energy. In Sec. LABEL:sec:normal_modes of the Supplemental Information (SI), the BO perturbation expansion is derived and expressed in normal mode coordinates. We then show in Section VI the extension to solids with practical formulae for implementation in first-principles software. In Section VII, we show that Allen’s energy is not the electron-phonon contribution to the total energy but a part of the phononic contribution to the total energy. We then recast in Section VIII Allen’s formula into the standard Allen-Heine-Cardona (AHC) [2, 3] theory for band renormalization thanks to an identically null term. In Section IX, we present a detailed analysis of the energy terms at the phonon zone-center, comparing a frozen-phonon approach and perturbation theory. We find excellent agreement for all terms, below 0.5% differences. We show size consistency by computing the diamond band renormalization using the Abinit code [28, 29, 71, 90] and an interpolation of the perturbed potential which includes dipole and quadrupole contributions [18, 17]. We show that the same results are obtained between the primitive and a supercell of diamond when no crystal symmetries are used.
@@
Finally, in Section X we compute the (fourth order) electron-phonon contribution to the total energy and approximate the effect of electron mass rescaling (sixth order), for two allotropes of carbon: the stable face-centered cubic diamond and the high-pressure metastable hexagonal diamond called lonsdaleite which is typically found in meteorites. For both materials, we report their converged total energy with momentum integration. We compute the electron-phonon forces and the new equilibrium lattice parameters of diamond and the internal parameter of lonsdaleite due to quasi-harmonic effects. At 0 K, we find that diamond is 52.7 meV/(2 atoms) more stable than lonsdaleite.
In addition to clarifying Allen’s work, the main findings of this manuscript are (i) an equation fulfilled by the exact total energy, Eq. (31), (ii) the derivation of all terms in a perturbation expansion series in terms of M-1/4 up to order 6 included, (iii) the identification of terms that do not come from the Born-Oppenheimer anharmonicity, including those coming from the vector potential, the electron-phonon coupling, mixed terms, and the renormalization of nuclei mass due to electronic inertial effects, and (iv) we derived a practical formula for the lowest-order electron-phonon contribution to the total energy, implement it and applied to study diamond and lonsdaleite.
@@ II Born-Oppenheimer-based approximations
We first consider finite systems and extend the theory to periodic crystals in Sec. VI. The total Hamiltonian of a system composed of electrons and nuclei can be expressed as [13, 14]
@@
where $\{{\mathbf{R}}\}$ denotes the set of nuclear positions, $\hat{T}^{\textrm{N}}$ is the nuclear kinetic operator, $\hat{P}_{\kappa\alpha}\equiv-i\partial/\partial R_{\kappa\alpha}$ is the canonical momentum operator for nuclei, with $M_{\kappa}$ the mass of nucleus $\kappa$ and $\alpha$ an index for cartesian axes. Electronic coordinates are implicit, unlike the nuclear ones. The global Hamiltonian in Eq. (1) is written in calligraphic style as well as the forthcoming associated total energies, which are independent of $\{{\mathbf{R}}\}$. Hartree atomic units are used throughout.
@@
The denominations “Born-Oppenheimer Hamiltonian”, “Born-Oppenheimer wavefunction” and “Born-Oppenheimer eigenenergy” for quantities in Eqs. (3) and (6) are well-established and consistent in the literature, but there is some confusion about the meaning of “Born-Oppenheimer approximation” and “adiabatic approximation” [7, 54, 34, 86, 81, 21]. In many publications, these are used interchangeably while in other they have different meanings. For example, in their 1972 publication, Ballhausen and Hansen [7] discuss various adiabatic approximations including “Born-Oppenheimer”, “Born-Huang”, “Herzberg-Teller”, and “crude”, based on wavefunctions of the type
[…]
that is, a simple product of a configuration-dependent (normalized) function of the nuclear configuration only $X_{i}\{{\mathbf{R}}\}$, with a single normalized electronic wavefunction, also possibly configuration-dependent. The electronic wavefunctions in these schemes change smoothly with the configuration. However, beyond the different “adiabatic approximations” used by Ballhausen and Hansen, the recent “exact factorization” [1, 81, 21] approach also has the form of Eq. (7) but is exact in that case. Therefore, it is not helpful to define the “adiabatic approximation” as being based on Eq. (7), which is too general.
[…]
@@
Figure 1: Example of coordinate diagram for the BO energy hypersurfaces $E^{\rm BO}_{j}\{{\mathbf{R}}\}$ of the electronic ground state ($j=0$) and first- and second-excited states ($j=1$ and $j=2$). Eigenenergies of the total Hamiltonian are also shown, labeled with the $i$ index. Unlike their wavefunctions, or the BO eigenenergies, they do not depend on the configuration of nuclei. The ground state of the total Hamiltonian has $i=0$ label. Usually, the corresponding wavefunctions will be made predominantly of the BO wavefunctions of one of the BO hypersurfaces, but there could be hybridized states. The chosen reference configuration is denoted by $\{\mathbf{R}^{\rm ref}\}$. In this example, we present a specific case where the $j=0$ BO hypersurface has two minima, a global one and a local one, that might both be chosen to be the reference state $\{\mathbf{R}_{j=0}^{0}\}$ in Sec. III. The $j=2$ BO hypersurface has one global minimum.
In the literature, there are several possible approximations to Eq. (8) in which only one term is kept. A first possible approximation consists in considering $E^{\rm BO}_{j}\{{\mathbf{R}}\}$ as an effective scalar potential for nuclear dynamics:
@@
A second possible approximation, still using one term in Eq. (8), which we refer to as the “optimal approximation”, is based on choosing $\chi_{ij}\{{\mathbf{R}}\}$ such that the expectation value of the Hamiltonian, Eq. (1), is stationary (minimal in the ground-state case). In that case, the nuclear wave function fulfills the following equation:
@@
where the projector on the $j$-BO manifold is defined by its matrix elements for two electronic and nuclear wavefunctions as follows,
[…]
or, in a more compact notation,
$|,\hat{\cal{P}}^{\textrm{BO}}_{j}\equiv\int\mathrm{d}\{{\mathbf{R}}\}\ket{\phi_{j}\{{\mathbf{R}}\}}\bra{\phi_{j}\{{\mathbf{R}}\}}$, (21)
while its orthogonal is
@@
The presence of the projector in Eq. (19) removes the phase arbitrariness that is present in Eq. (14). In the literature dealing with phonon chirality [11], a simplified version of Eq. (15) is often encountered where the vector potential is considered but not the Born term, see Sec. LABEL:sec:hybrid of SI.
Both approximations, Eqs. (11) and (15), to the exact Schrödinger equation, Eq. (1), have been referred to as “adiabatic approximation” and/or as “Born-Oppenheimer approximation” in the literature, with the following approximate total wavefunctions:
@@
The introduction of the reference set of BO wavefunctions, and its use in Eqs. (8), (23a), and (23b), already allows one to obtain an exact representation of the total wavefunction, as well as to introduce different adiabatic approximations, but even yields a nuclear wavefunction equation that must be fulfilled by the exact eigenenergy.
@@
with the term between parentheses defining a pseudo-inverse, limited to that subspace. Although Green’s function might not be defined for specific real $z$ values above the ground-state energy $z=\mathcal{E}_{0}$, the restriction to the space orthogonal to the BO-wavefunctions removes this problem up to $z$ being equal to the lowest excited-state BO eigenenergy. We also assume that $z=\mathcal{E}_{i}$ is a value for which the projected Green’s function $\hat{\mathcal{G}}_{\perp j}(z,\{{\mathbf{R}}\})$ exists.
@@
is an operator similar to self-energy operators found in the context of many-body perturbation theory [48].
Importantly, Eq. (31) is an exact version of the approximate Eqs. (11) and (15), and focuses only on nuclear dynamics where all electronic degrees of freedom are condensed in the self-energy operator. It is a key result of this work.
The Green’s function of the full Hamiltonian fulfills a Dyson-type equation: